OpenStudy (anonymous):

For what values of A and B is the polynomial \[f(x,y)= Ax ^{4}+6x ^{2}y ^{2}+By ^{4}\] function harmonic in the whole xy-plane?

OpenStudy (anonymous):

\[df/dx = 4Ax ^{3}+12xy ^{2}\] \[d^{2}f /dx ^{2}=12Ax ^{2}+12y ^{2}\] \[df/dy=12x ^{2}y+4By ^{3}\] \[d ^{2}f/dx ^{2}=12x ^{2}+12By ^{2}\] the condition for harmonic function is: \[d ^{2}f/dx ^{2}+d ^{2}f/dy ^{2}=0\] so A=-1 and B=-1

OpenStudy (anonymous):

@Wislar

OpenStudy (anonymous):

So... \[12Ax ^{2}+12y ^{2}+12x ^{2}+12By ^{2}=0\] \[12A+12=0 \\ 12B+12=0\] Right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Thanks!